OpenStudy (anonymous):

What is the ratio between the 12th and 11th terms in the Fibonacci Sequence? Round your answer to five decimal places. Term 1 2 3 4 5 6 7 8 9 Fibonacci Number 1 1 2 3 5 8 13 21 34

OpenStudy (anonymous):

can u expllain is it a fration

OpenStudy (anonymous):

next numbers are 55 89 144

OpenStudy (anonymous):

89 = 11th and 144 = 12th ratio you require is 144/89 as a decimal this is 1.61798 to 5 dec places

OpenStudy (anonymous):

but how is that a ratio

OpenStudy (anonymous):

This is also known as the golden ratio.

OpenStudy (anonymous):

1.61798 : 1 \[\frac{1.61798}{1}\] 1.61798 It is standard practice to omit the one when the context makes the meaning clear.

OpenStudy (anonymous):

whatever it is, it will be close to \[\phi =\frac{1+\sqrt{5}}{2}\]